OpenStudy (anonymous):

can someone show me the steps to solve x/6=x/7+5

OpenStudy (anonymous):

multiply with 6 and 7 than bring x to one side. voila

OpenStudy (anonymous):

\[\frac{ x }{ 6 }=\frac{ x }{ 7 }+5\] to better understand here is the actual problem in the correct equation

OpenStudy (anonymous):

Move the x's to one side. Make the denominators equal. Minus the two terms, and move the denominator over to the over side and multiply 5 by the denominator (42) and then divide by the number next to x on the side where 5(42) is. and voila

OpenStudy (anonymous):

7x/42 - 6x/42 = 5 x/42=5 x=5(42)

OpenStudy (calculusfunctions):

\[\frac{ x }{ 6 }=\frac{ x }{ 7 }+5\] Multiply both sides by 42 (least common multiple of 6 and 7), to obtain 7x = 6x + 210 ∴ x = 210 Do you understand?

OpenStudy (anonymous):

so step 2 would look like x/7-x/6=5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then you make the denominators the same

OpenStudy (anonymous):

which would be 42,

OpenStudy (anonymous):

yes since 6 X 7 = 42

OpenStudy (anonymous):

this is the part i get confused at

OpenStudy (anonymous):

would it look like x/42?

OpenStudy (anonymous):

but in order to make the denominators equation, you need to times the numerator by the amount that the denominator needs to make 42. eg. 6 X ? = 42 (7) so times the top x by 7 which becomes 7x. and the other becomes 6x since you need to times 7 by 6 to make 42. so it would be 7x/42 - 6x/42 7x-6x = x so it will eventually become x/42

OpenStudy (anonymous):

ahhh ok

OpenStudy (anonymous):

so then is x/42=5

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

and then simplify from there

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

how do i get x by itself?

OpenStudy (anonymous):

move the 42 over to the other side by timesing both sides by 42.

OpenStudy (anonymous):